Given the demand function, P36 Q², and the supply function PS-30+ 8. Assuming pure competition, what is the consumers' surplus?


O a. 62.33


b. 32.33


C. 52.33


6.42.33


1
Expert's answer
2021-06-28T18:17:03-0400

"p=36-Q ^2\\\\\n \nQ= \\sqrt{-p+36}\\\\\n\u200b\n \np=30+8Q\\\\\nQ= \\frac{p\u221230}{8}\\\\\n\u200b\n \nQs=Qd\\\\\n \\sqrt{-p+36}=\\frac{p\u221230}{8}\\\\\np=35.523\\\\Q= \\frac{p\u221230}{8}\\\\Q=0.690\n\u200b"


"0= \\sqrt{-p+36} \\\\\np=36\n\u200b"

"\frac{1}{2}\times 0.690\times(36-35.523)=0.164565"